The length of a rectangle is 2ft less than three times the width of the triangle. The area of the rectangle is 65ft2. Find the length and width of the rectangle.
Answer by checkley71(8403)

L=3W-2 I'M ASSUMING YOU MEAN THE WIDTH OF THE RECTANGLE NOT THE TRIANGLE(?).
L*W=AREA
(3W-2)W=65
3W^2-2W-65=0
(3W+13)(W-5)=0
W-5=0
W=5 FOR THE WIDTH.
L=3*5-2=15-2=13 FOR THE LENGTH.
PROOF:
5*13=65
65=65